OpenStudy (anonymous):

What is the value of z in the equation 2(4z –6 –6) = 170 –46? 12.5 15.5 16.5 18.5

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

combine like terms first

OpenStudy (misty1212):

then divide both sides by 2

OpenStudy (anonymous):

hi can you help!

OpenStudy (misty1212):

\[2(4z-12)=124\] is a start

OpenStudy (misty1212):

that is what you get when you do the arithmetic, \(-6-6=-12\) and \(170-46=124\)

OpenStudy (misty1212):

then divide by 2 and you get \[4z-12=124\] two more steps

OpenStudy (misty1212):

a) add \(12\) to both sides, then b) divide both sides by \(4\)

OpenStudy (anonymous):

What is the value of z in the equation 2(4z –6 –6) = 170 –46? 12.5 15.5 16.5 18.5 evaluate the expression first 2(4z-6-6)=170-46 2(4z-12)=170-46 8z-24=124 8z=124+24 8z=148 Divide both sides by 8...we come up this z=18.5

OpenStudy (misty1212):

or you can just read the answer above

OpenStudy (anonymous):

Thus, z=18.5

OpenStudy (misty1212):

or you can click on this link http://www.wolframalpha.com/input/?i=+2%284z+%E2%80%936+%E2%80%936%29+%3D+170+%E2%80%9346

OpenStudy (anonymous):

Any question bothering to you?

OpenStudy (anonymous):

@agl202301 : if their is somewhat to ask just free to post.